OpenStudy (anonymous):

Let A and B be events in a sample space S such that P(A) = 0.6, P(B) = 0.4, and P(A intersection B) = 0.18. Find the probabilities below. (a) P(A|B) (b) P(B|A)

OpenStudy (anonymous):

I got B

OpenStudy (anonymous):

its to separate problems

OpenStudy (anonymous):

oh I see i'll still try to solve it

OpenStudy (anonymous):

ok first you do experimental probability by flipping a coin then you can write out the outcomes and divide them and solve

OpenStudy (anonymous):

I hope it helped good luck!

OpenStudy (bradely):

P(A/B) =P(A intersection B )/P(B) =0.18/0.4 P(B/A) =P(A intersection B )/P(A) =018/0.6 Source: http://www.mathskey.com

OpenStudy (anonymous):

thanks bradely.that helped me a lot.